# Boosting Generic Semi-Supervised Medical Image Segmentation via Diverse Teaching and Label Propagation

### v1

@@ III-B Pseudo Labeling with Diverse Teaching
For the unlabeled data flow, i.e., $\mathcal{D}(x^{u};\theta)$ as decoder, only takes $x_{u}$ as input to obtain the multi-scale features $h^{u}_{i}$. To further train the $\mathcal{D}(x^{u};\theta)$ branch, we design Diverse Teaching strategy to generate high quality Pseudo Labels.
@@
Unsupervised Training with $\mathbf{D}(x^{u};\theta)$
@@
$\mathcal{L}_{u}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\mathcal{L}_{DiceCE}(p^{u;\theta}_{i},y^{\xi,\psi})$ (11)
@@ III-C Global-local consistency learning
To fully explore the data structure, we further adopt the cross-set CutMix strategy [34] and masked image modeling [35, 36] to achieve global-local consistency learning, thus improving the model generalization and reducing domain shifts.
@@
$\mathcal{L}_{mix}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\mathcal{L}_{DiceCE}(p^{u;\theta}_{i,mix},y^{t}_{i,mix}),$ (13)
@@
III-D Masked Reconstruction and Knowledge Distillation
The hard pseudo labels generated by the dual teachers still inevitably contain noise, we further try to exploit the soft predictions generated by the mean-teacher Encoder $\mathcal{E}(x^{l},x^{u};\xi^{T})$ and Decoder $\mathcal{D}(x^{u};\theta^{T})$ to distill knowledge from teacher to the masked student, thereby capturing the contextual of the medical image without the noise of the hard pseudo labels. To this end, we minimize the difference between the voxel-wise logits generated by the teacher network given the original input $\bm{x}_{i}$ and the voxel-wise logits produced by the student using the masked input $\bm{x}_{i,mask}$. The objective function is expressed as,
$\mathcal{L}_{rec}=\frac{1}{N_{L}+N_{U}}\sum_{i=0}^{N_{L}+N_{U}}\frac{\|\ p^{u,l;\theta}_{i,mask}-p^{u,l;\theta^{t}}_{i}\|\ _{2}^{2}}{\|\ p^{u,l;\theta^{t}}_{i}\|\ _{2}^{2}}.$ (16)
To furthermore reduce the potential noise in hard pseudo labels generated via eq. 10, for the unlabeled data, we distill knowledge from the Decoder branches $\mathcal{D}(x^{l};\psi)$ and $\mathcal{D}(x^{l},x^{u};\xi)$ to the $\mathcal{D}(x^{u};\theta)$ decoder, respectively. This is accomplished by minimizing the soft dice loss $\mathcal{L}_{sdice}$ [38] using the predictions of $p^{u;\xi}$ and $p^{u;\psi}$ .
$\mathcal{L}_{kd}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\left(\mathcal{L}_{sdice}\left(p^{u;\theta}_{i},p^{u;\xi}\right)+\mathcal{L}_{sdice}\left(p^{u;\theta}_{i},p^{u;\psi}\right)\right)$ (17)
@@ III-E Voxel-level Label Propagation
The global-local consistency learning can explore the inter-sample and intra-sample data structure. However, the pairwise similarities and consequently the voxel-level remain to be solved. In this section, we propose the voxel-level propagation strategy to fully explore the potential of unlabeled data. First, we extract features $e_{1}$ and $e_{2}\in\mathbb{R}^{D\times LHW}$ through linear layers after the encoder of the network, where $D$ is the channel dimension and $LHW$ is the number of feature vectors.
@@
$p_{i,coor}^{u;\theta}=\mathcal{D}(\mathcal{E}(x^{u}_{i};\xi);\theta)\cdot\mathcal{C},$ (19)
The resulting $\mathbf{z}_{i}^{u}$ emphasizes the pairwise similarities of the same object through the correlation map.
Therefore, a correlation loss $\mathcal{L}_{coor}^{u}$ can be calculated between $\mathbf{z}_{i}^{u}$ and the pseudo labels $y_{i}^{t}$ as the supervision, which can be written as follows:
$\mathcal{L}_{coor}^{u}=\frac{1}{N_{U}}\sum^{N_{U}}_{i=1}(\mathcal{L}_{DiceCE}(p_{i,coor}^{u;\theta},y_{i}^{t})).$ (20)
Similarly, for the labeled images branch $\mathcal{D}(x_{i}^{l};\xi)$, we also compute the $\mathcal{L}_{coor}^{l}$. Then, the correlation loss is obtained via:
$\mathcal{L}_{coor}=\mathcal{L}_{coor}^{u}+\mathcal{L}_{coor}^{l}.$ (21)
@@ III-F Total Taring loss
Integrating these objectives introduced in Eqs. 2, 3, 12,  15,  17,  16,  21 together, the final loss function as follows:
$\mathcal{L}=\mathcal{L}_{deno}+\mathcal{L}_{diff}+\mathcal{L}_{mix}+\alpha\mathcal{L}_{mic}+\beta\mathcal{L}_{kd}+\gamma\mathcal{L}_{rec}+\eta\mathcal{L}_{crro}.$ (22)
@@ IV-A Datasets and Implementation Details
Our method is evaluated on five publicly accessible datasets, including three semi-supervised benchmark datasets, i.e., the LA dataset [39], Synapse dataset [40], and AMOS dataset [41], one dataset MMWHS dataset [42] for UMDA, and two datasets M&Ms [43] for Semi-MDG. To assess the network’s prediction, the Dice metric and the average surface distance (ASD) are employed. In the context of SSMIS tasks, following previous works [2], the additional Jaccard and HD95 metrics are also utilized.
@@ IV-B Competitors
General SSMIS methods : UA-MT [17]; URPC [44]; CPS [45]; MC-Net [46]; $S$-Ne [47]; DePL [48]; BCP [18]; OTCMC [19]; GapMatch [22]; w2sPC [21]. Imbalance SSMIS methods: Adsh [49] ; SimiS [50]; DHC [51]; AllSpark [52] ; A&D [10]; SKCDF [53].
@@
PnP-AdaNet [25];AdaOutput [6] , CycleGAN [4], CyCADA [23] , SIFA [26], DSAN [28], LMISA-3D [54], DDSP [5], A&D [10].
@@
SDNet+Aug [31] , LDDG [55], SAML [56], DGNet [30] , vMFNet [8] , Meta [57], StyleMatch [58] , EPL [59]. A&D [10].
TABLE II: Comparisons on the LA dataset with 5% labeled data.
@@
w2sPC [21] (MedIA’25) 89.02 79.83 10.23 2.18
GapMatch [22] (AAAI’25) [22] 88.3 - - 1.77
OTCMC[19] (TMI’25) 87.97 78.65 8.39 2.02
[…]
TABLE III: Comparisons on the LA dataset with 10% labeled data.
Method Scans used Metrics
@@
w2sPC [21] (MedIA’25) 90.23 81.52 7.16 1.95
GapMatch [22] (AAAI’25) 91.0 - - 1.46
OTCMC[19] (TMI’25) 90.26 82.34 5.96 1.65
@@
TABLE V: Quantitative comparisons on 10% labeled Synapse dataset.
@@
TABLE VI: RESULTS ON AMOS DATASET FOR SSMIS TASK.
@@
TABLE VII: Results on two settings, i.e., MR to CT and CT to MR, of MMWHS dataset for UMDA task.
@@
TABLE VIII: Results on 2% and 5% labeled data settings of M&Ms dataset for Semi-MDG task.
@@
Figure 2: Visualization results of different semi-supervised segmentation techniques are depicted, utilizing 5% labeled data alongside ground truth on the LA dataset.
Figure 3: Visualization results of different semi-supervised segmentation techniques are depicted, utilizing 20% labeled data alongside ground truth on the Synapse dataset.
@@ IV-C Comparison with state-of-the-art methods
Tab. II, Tab. III, Tab. IV, Tab. V, and Tab. VI show quantitative comparison results for the LA, Synapse, and AMOS datasets. Our proposed DTLP-Net achieves notable enhancements in all evaluation metrics by a significant margin in different training scenarios, effectively leveraging the potential of unlabeled data. Specifically, on the LA dataset, our method achieves SOTA segmentation performance, surpassing the fully supervised method by 0.16% (91.63% vs. 91.47%) and 0.96%(92.43% vs. 91.47%) in dice performance with only 5% and 10% labeled data, respectively. Our method achieves notable enhancements in Dice, Jaccard, 95HD, and ASD metrics, surpassing the second-best A&D [10] performance by 1.7%, 2.77%, 0.85, and 0.45, respectively, for the 5% setting. Similarly, for the 10% data, we outperform the second-best approach GapMatch [22] 1.43% and 0.16% in Dice and ASD performance, respectively. These results are obtained without conducting any post-processing, ensuring fair comparisons with other methods.
Notably, on the Synapse dataset, our method achieves SOTA segmentation performance on most types of organs, which shows the promising ability to solve traditional SSMIS tasks, as shown in Tabs. IV and V . Furthermore, by incorporating the class-imbalanced designs from [10], our method successfully segments the minor classes RAG and LAG, resulting in an improved overall Dice score of 66.01% under the 20% labeled setting. This achievement demonstrates a significant improvement (5.13% in Dice) compared to previous methods.
We further evaluate our method on the AMOS dataset, and the experimental results are shown in Tab. VI. Similarly, our method also achieves SOTA performance on the AMOS dataset. Specifically, as shown in Tab. VI, on the AMOS dataset with 2% labeled data, our method outperforms the second-best method, i.e., AllSpark [52] with 8.79% in Dice and 4.07 in ASD.
@@
Table VII presents the outcomes of our method for the UMDA task on the MMWHS dataset. In the Dice metric on MR to CT task, our methods outperform the fully-supervised method by 1.2% (92.1% vs. 90.9%). Moreover, from CT to MR task, our methods are merely 0.2% inferior to the fully-supervised method, showing its extraordinary performance. Compared to UMDA methods that adopt image-level (CycleGAN [4]), feature-level (DSAN [28]) or both (CyCADA [23], SIFA [26], DDSP [5] ) alignments to mitigate domain shifts, our method achieves comparable performance, surpassing the second-best DDSP [5] performance by 3.1% (92.1% vs. 89.0%) and 2.5% (83.4% vs. 81.9%) in Dice on two UMDA tasks. Furthermore, when compared with generic semi-supervised segmentation method, e.g., A&D [10], our method outperform 12.0% (83.4% vs. 71.4%) in Dice on CT to MR task.
@@
Tab. VIII shows the comparison results of Semi-MDG methods on M&Ms dataset. Compared to EPL [59], which adopts Fourier transformation [59] to deal with the domain shift, our methods achieve an improvement of 0.96% (85.46% vs. 84.52%) and 0.5% (86.22% vs. 85.72) with 2% and 5% labeled data, respectively. Notably, when compared with other SOTA pure semi-supervised segmentation methods, e.g., BCP [1], our method shows solid performance gains of 9.89% (85.46% vs. 77.94%) and 9.17%(86.22% vs. 77.05%), respectively. Furthermore, when compared with other SOTA pure semi-supervised segmentation methods, e.g., BCP [1], our method shows solid performance gains. Furthermore, we surpass the generic semi-supervised segmentation method A&D [10] by 4.15% (85.46% vs. 81.31%) and 2.91% (86.22% vs. 83.31%), respectively. By leveraging the global - local data structure and devising an effective teaching strategy, our method generates more reliable pseudo - labels and mitigates the domain shift.
TABLE IX: Results of the ablation experiments carried out on the LA dataset, employing 5% labeled and 10% data.
@@
TABLE X: Results for ablation experiments which are conducted on the Synapse dataset using 20 % labeled data.
@@ IV-D1 Effects of different components
To validate the effectiveness of each component of our method in Eq. 22, i.e., global-local consistency loss $\mathcal{L}_{mix}$ and $\mathcal{L}_{mic}$, knowledge distillation loss $\mathcal{L}_{kd}$, Masked image Reconstruction $\mathcal{L}_{krec}$, and voxel-level label Propagation loss $\mathcal{L}_{corr}^{u}$, we conduct ablation studies on the LA dataset across two distinct semi-supervised configurations and Synapse dataset with 20% labeled data for training. The outcomes of these ablation experiments are meticulously documented in Table IX and X, respectively.
@@
To fully explore the data structure, except for cross-set CutMix Strategy that engenders novel training samples that fill the void between input samples to regularize the global distributional smoothness [2], we further adopt masking image modeling to learn the local semantics of the data. As illustrated in Tab. IX and X, it is apparent that GLCL results in improved model performance. For instance, from Table IX, by further adopting masking image modeling, the Dice performance improves by 0.61% (90.94 vs. 90.35%) and 0.65% (92.27% vs. 91.61%) with 5% and 10% labeled data for training, respectively. Both cross-set CutMix and masking image modeling can reduce the domain shifts and improve the data-level diversity, thus improving the model generalization ability. Similar conclusions can be drawn from Table X. However, compared with the LA dataset, the Synapse dataset challenges with class imbalance and contains many small foregrounds; random masking may mask out the foregrounds, leading to suboptimal performance.
@@
To further mitigate the potential noise in the hard pseudo-labels generated through Equation 10, we perform knowledge distillation from the Decoder branches $\mathcal{D}(x^{l};\psi)$ and $\mathcal{D}(x^{l},x^{u};\xi)$ to the $\mathcal{D}(x^{u};\theta)$ decoder, respectively. As can be clearly observed from Tables IX and X, knowledge distillation is more effective on the Synapse dataset. However, it exhibits relatively minor improvements on the LA dataset. The underlying reason is that the Synapse dataset exhibits class imbalance and numerous small foregrounds. This results in the hard pseudo-labels prior being more prone to noise. However, distilling soft predictions from the other two branches may mitigate this issue somewhat. In addition, the reconstruction of the features of the masked input volumes promotes contextual consistency, thereby enhancing the segmentation performance. As presented in Tables IX, through the additional implementation of masked reconstruction, the Dice performance increases from 90.73% to 91.16%.
@@
To comprehensively explore the potential of unlabeled data, we investigate the label propagation strategy to capture the voxel-level pairwise similarities, thereby improving the segmentation performance. As can be observed from Tables IX and X, the Dice performance can be significantly enhanced.

### v3

@@ III-B Pseudo Labeling with Diverse Teaching
For the unlabeled data flow, i.e., $\mathcal{D}(x^{u};\theta)$ as decoder, only takes $x_{u}$ as input to obtain the multi-scale features $h^{u}_{i}$. To further train the $\mathcal{D}(x^{u};\theta)$ branch, we design Diverse Teaching strategy to generate high quality pseudo labels, as illustrated in Fig. 1 (b).
@@
Unsupervised Training with $\mathcal{D}(x^{u};\theta)$
@@
$\small\mathcal{L}_{u}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\mathcal{L}_{DiceCE}(p^{u;\theta}_{i},y^{\xi,\psi})$ (7)
@@ III-C Global-local consistency learning
To fully explore the data structure, we further adopt the cross-set CutMix strategy [30] and masked image modeling [31] to achieve global-local consistency learning, thus improving the model generalization and reducing domain shifts.
@@
$\small\mathcal{L}_{mix}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\mathcal{L}_{DiceCE}(p^{u;\theta}_{i,mix},y^{t}_{i,mix}),$ (9)
@@
III-D Knowledge Distillation and Masked Reconstruction
The hard pseudo labels generated by the dual teachers via eq. 6 still inevitably contain noise, especially in the scenario where there is class imbalance. Therefore, for the unlabeled data, we distill knowledge from the Decoder branches $\mathcal{D}(x^{l};\psi)$ and $\mathcal{D}(x^{l},x^{u};\xi)$ to the student $\mathcal{D}(x^{u};\theta)$ decoder, respectively. This is accomplished by minimizing the soft dice loss $\mathcal{L}_{sdice}$ [33] using the predictions of $p^{u;\xi}$ and $p^{u;\psi}$ .
$\small\mathcal{L}_{kd}=\frac{1}{N_{U}}\sum_{i=0}^{N_{U}}\left(\mathcal{L}_{sdice}\left(p^{u;\theta}_{i},p^{u;\xi}\right)+\mathcal{L}_{sdice}\left(p^{u;\theta}_{i},p^{u;\psi}\right)\right).$ (12)
In addition, we further try to exploit the soft predictions generated by the mean-teacher Encoder $\mathcal{E}(x^{l},x^{u};\xi^{T})$ and Decoder $\mathcal{D}(x^{u};\theta^{T})$ to reconstruct the masked student, thereby capturing the contextual of the medical image with less noise. Specifically, we further minimize the difference between the voxel-wise logits generated by the second teacher network given the original input $\bm{x}_{i}$ and the voxel-wise logits produced by the student using the masked input $\bm{x}_{i,mask}$. The objective function is expressed as,
$\small\mathcal{L}_{rec}=\frac{1}{N_{L}+N_{U}}\sum_{i=0}^{N_{L}+N_{U}}\frac{\|\ p^{u,l;\theta}_{i,mask}-p^{u,l;\theta^{t}}_{i}\|\ _{2}^{2}}{\|\ p^{u,l;\theta^{t}}_{i}\|\ _{2}^{2}}.$ (13)
@@ III-E Voxel-level Label Propagation
The global-local consistency learning can explore the inter-sample and intra-sample data structure. However, the pairwise similarities on the voxel-level remain to be solved. In this section, we propose the voxel-level propagation strategy to fully explore the potential of unlabeled data. First, we extract features $e_{1}$ and $e_{2}\in\mathbb{R}^{D\times LHW}$ through linear layers after the encoder of the network, where $D$ is the channel dimension and $LHW$ is the number of feature vectors.
@@
$p_{i,corr}^{u;\theta}=\mathcal{D}(\mathcal{E}(x^{u}_{i};\xi);\theta)\cdot\mathcal{C},$ (15)
the resulting $p_{i,corr}^{u;\theta}$ emphasizes the pairwise similarities of the same object through the correlation map. Then, a correlation loss $\mathcal{L}_{corr}^{u}$ can be calculated as follows:
$\mathcal{L}_{corr}^{u}=\frac{1}{N_{U}}\sum^{N_{U}}_{i=1}(\mathcal{L}_{DiceCE}(p_{i,corr}^{u;\theta},y_{i}^{t})).$ (16)
Similarly, for the labeled images branch $\mathcal{D}(x_{i}^{l};\xi)$, we also compute the $\mathcal{L}_{corr}^{l}$. Then, the correlation loss is obtained via:
$\mathcal{L}_{corr}=\mathcal{L}_{corr}^{u}+\mathcal{L}_{corr}^{l}.$ (17)
@@ III-F Total Taring loss
Integrating these objectives introduced in Eqs. 2, 3, 8,  11,  12,  13,  17 together, the final loss function as follows:
$\small\mathcal{L}=\mathcal{L}_{diff}+\mathcal{L}_{sup}+\mathcal{L}_{mix}+\alpha\mathcal{L}_{mic}+\beta\mathcal{L}_{kd}+\gamma\mathcal{L}_{rec}+\eta\mathcal{L}_{corr},$ (18)
where $\alpha$, $\beta$, $\gamma$, and $\eta$ are the parameters to control the importance of the loss.
@@ IV-A Datasets and Implementation Details
Our method is evaluated on five publicly accessible datasets, including three semi-supervised benchmark datasets, i.e., the LA dataset [34], Synapse dataset [35], and AMOS dataset [36], one dataset MMWHS [37] for UMDA, and one dataset M&Ms [38] for Semi-MDG. To assess the network’s prediction, the Dice metric and the average surface distance (ASD) are employed. In the context of SSMIS tasks, following previous works [2], the additional Jaccard and HD95 metrics are also utilized.
@@ IV-B Competitors
General SSMIS methods : UA-MT [17]; URPC [39]; CPS [40];DTC [41]; SASSNet [42]; MC-Net [43]; $S$-Ne [44]; DePL [45]; BCP [1]; OTCMC [18]; GapMatch [20]; w2sPC [19]. Imbalance SSMIS methods: Adsh [46] ; SimiS [47]; DHC [48]; AllSpark [49] ; A&D [12]; InterTeach [4]; SKCDF [50].
@@
PnP-AdaNet [23], CycleGAN [6], CyCADA [21] , DSAN [24], LMISA-3D [51], DDSP [7].
@@
SDNet+Aug [27] , LDDG [52], SAML [53], DGNet [26] , vMFNet [10], Meta [54], StyleMatch [55], EPL [56].
TABLE II: Comparisons on the LA dataset with 5% and 10% labeled data, respectively.
@@
InterTeach [4] 89.76 81.49 6.31 1.75
w2sPC [19] 89.02 79.83 10.23 2.18
GapMatch [20] 88.3 - - 1.77
OTCMC[18] 87.97 78.65 8.39 2.02
@@
InterTeach [4] 91.11 83.72 5.19 1.55
w2sPC [19] 90.23 81.52 7.16 1.95
GapMatch [20] 91.0 - - 1.46
OTCMC[18] 90.26 82.34 5.96 1.65
@@
TABLE IV: Quantitative comparisons on 10% labeled Synapse dataset.
@@
TABLE V: RESULTS ON AMOS DATASET FOR SSMIS TASK.
@@
Figure 2: Visualization results with 5% labeled data on the LA dataset.
TABLE VI: Results on two settings, i.e., MR to CT and CT to MR, of MMWHS dataset for UMDA task.
@@
TABLE VII: Results on 2% and 5% labeled data settings of M&Ms dataset for Semi-MDG task.
@@ IV-C Comparison with state-of-the-art methods
Tab. II, Tab. III, Tab. IV, and Tab. V show quantitative comparison results for the LA, Synapse, and AMOS datasets. Our proposed DTLP-Net achieves notable enhancements in all evaluation metrics by a significant margin in different training scenarios, effectively leveraging the potential of unlabeled data. Specifically, on the LA dataset, our method achieves SOTA segmentation performance, surpassing the fully supervised method by 0.16% (91.63% vs. 91.47%) and 0.96%(92.43% vs. 91.47%) in dice performance with only 5% and 10% labeled data, respectively. Our method achieves notable enhancements in Dice, Jaccard, 95HD, and ASD metrics, surpassing the second-best A&D [12] performance by 1.7%, 2.77%, 0.85, and 0.45, respectively, for the 5% setting. Similarly, with 10% labeled data for training, we outperform the second-best approach GapMatch [20] by 1.43% and 0.16 in Dice and ASD performance, respectively. These results are obtained without conducting any post-processing, ensuring fair comparisons with other methods.
Notably, on the Synapse dataset, our method achieves SOTA segmentation performance on most types of organs, which shows the promising ability to solve traditional SSMIS tasks, as shown in Tabs. III and IV . Furthermore, by incorporating the class-imbalanced designs from [12], our method successfully segments the minor classes RAG and LAG, resulting in an improved overall Dice score of 66.01% under the 20% labeled setting. This achievement demonstrates a significant improvement (5.13% in Dice) compared to previous methods.
We further evaluate our method on the AMOS dataset, and the experimental results are shown in Tab. V. Similarly, our method also achieves SOTA performance on the AMOS dataset. Specifically, as shown in Tab. V, on the AMOS dataset with 2% labeled data, our method outperforms the second-best method, i.e., AllSpark [49] with 8.79% in Dice and 4.07 in ASD.
@@
Figure 3: Visualization results with 20% labeled on the Synapse dataset.
@@
Table VI presents the outcomes of our method for the UMDA task on the MMWHS dataset. In the Dice metric on MR to CT task, our methods outperform the fully-supervised method by 1.2% (92.1% vs. 90.9%). Moreover, from CT to MR task, our methods are merely 0.2% inferior to the fully-supervised method, showing its extraordinary performance. Compared to UMDA methods that adopt image-level (CycleGAN [6]), feature-level (DSAN [24]) or both (CyCADA [21], DDSP [7] ) alignments to mitigate domain shifts, our method achieves comparable performance, surpassing the second-best DDSP [7] performance by 3.1% (92.1% vs. 89.0%) and 1.5% (83.4% vs. 81.9%) in Dice on two UMDA tasks. Furthermore, when compared with the generic semi-supervised segmentation method, e.g., A&D [12], our method outperforms by 12.0% (83.4% vs. 71.4%) in Dice on the CT to MR task.
@@
Tab. VII shows the comparison results of Semi-MDG methods on M&Ms dataset. Compared to EPL [56], which adopts Fourier transformation [56] to deal with the domain shift, our methods achieve an improvement of 0.94% (85.46% vs. 84.52%) and 0.5% (86.22% vs. 85.72) with 2% and 5% labeled data, respectively. Notably, when compared with other SOTA pure semi-supervised segmentation methods, e.g., BCP [1], our method shows solid performance gains of 9.81% (85.46% vs. 75.65%) and 9.17%(86.22% vs. 77.05%), respectively. Furthermore, we surpass the generic semi-supervised segmentation method A&D [12] by 4.15% (85.46% vs. 81.31%) and 2.91% (86.22% vs. 83.31%), respectively. By leveraging the global - local data structure and devising an effective teaching strategy, our method generates more reliable pseudo - labels and mitigates the domain shift.
TABLE VIII: Results of the ablation experiments carried out on the LA dataset, employing 5% labeled and 10% data.
@@
TABLE IX: Results for ablation experiments which are conducted on the Synapse dataset using 20 % labeled data.
@@ IV-D1 Effects of different components
To validate the effectiveness of each component of our method in Eq. 18, i.e., global-local consistency loss $\mathcal{L}_{mix}$ and $\mathcal{L}_{mic}$, knowledge distillation loss $\mathcal{L}_{kd}$, masked image reconstruction $\mathcal{L}_{rec}$, and voxel-level label propagation loss $\mathcal{L}_{corr}^{u}$, we conduct ablation studies on the LA dataset across two distinct semi-supervised configurations and Synapse dataset with 20% labeled data for training. The outcomes of these ablation experiments are meticulously documented in Table VIII and IX, respectively.
@@
To fully explore the data structure, except for cross-set CutMix Strategy that engenders novel training samples that fill the void between input samples to regularize the global distributional smoothness [2], we further adopt masking image modeling to learn the local semantics of the data. As illustrated in Tab. VIII and IX, it is apparent that GLCL results in improved model performance. For instance, from Table VIII, by further adopting masking image modeling, the Dice performance improves by 0.59% (90.94 vs. 90.35%) and 0.66% (92.27% vs. 91.61%) with 5% and 10% labeled data for training, respectively. Both cross-set CutMix and masking image modeling can reduce the domain shifts and improve the data-level diversity, thus improving the model generalization ability. Similar conclusions can be drawn from Table IX. However, compared with the LA dataset, the Synapse dataset challenges with class imbalance and contains many small foregrounds, random masking may mask out the foregrounds, leading to suboptimal performance.
@@
To further mitigate the potential noise in the hard pseudo-labels generated through Eq. 6, we perform knowledge distillation from the Decoder branches $\mathcal{D}(x^{l};\psi)$ and $\mathcal{D}(x^{l},x^{u};\xi)$ to the $\mathcal{D}(x^{u};\theta)$ decoder, respectively. As can be clearly observed from Tables VIII and IX, knowledge distillation is more effective on the Synapse dataset. However, it exhibits relatively minor improvements on the LA dataset. The underlying reason is that the Synapse dataset exhibits class imbalance and numerous small foregrounds. This results in the hard pseudo-labels prior being more prone to noise. However, distilling soft predictions from the other two branches may mitigate this issue somewhat. In addition, the reconstruction of the features of the masked input volumes promotes contextual consistency, thereby enhancing the segmentation performance. As presented in Tables VIII, through the additional implementation of masked reconstruction, the Dice performance increases from 90.73% to 91.16%.
@@
To comprehensively explore the potential of unlabeled data, we investigate the label propagation strategy to capture the voxel-level pairwise similarities, thereby improving the segmentation performance. As can be observed from Tables VIII and IX, the Dice performance can be significantly enhanced.